The potential energy of a satellite orbiting the earth is-X joule the energy required to throw it out of the earth's gravitational field is



1
Expert's answer
2021-07-22T10:03:15-0400
"U=-\frac{GmM_E}{(R+h)}"

"\Delta U=\frac{GmM_E}{(R)}-\frac{GmM_E}{(R+h)}"